# Balanced Rate-Distortion Optimization in Learned Image Compression

### v1

@@ Abstract
Learned image compression (LIC) using deep learning architectures has seen significant advancements, yet standard rate-distortion (R-D) optimization often encounters imbalanced updates due to diverse gradients of the rate and distortion objectives. This imbalance can lead to suboptimal optimization, where one objective dominates, thereby reducing overall compression efficiency. To address this challenge, we reformulate R-D optimization as a multi-objective optimization (MOO) problem and introduce two balanced R-D optimization strategies that adaptively adjust gradient updates to achieve more equitable improvements in both rate and distortion. The first proposed strategy utilizes a coarse-to-fine gradient descent approach along standard R-D optimization trajectories, making it particularly suitable for training LIC models from scratch. The second proposed strategy analytically addresses the reformulated optimization as a quadratic programming problem with an equality constraint, which is ideal for fine-tuning existing models. Experimental results demonstrate that both proposed methods enhance the R-D performance of LIC models, achieving around a 2% BD-Rate reduction with acceptable additional training cost, leading to a more balanced and efficient optimization process. The code will be made publicly available.
@@ 3.1 Preliminaries: rate-distortion optimization in learned image compression
  • Encoding and Quantization: First, each data point $x$ is mapped to a de-correlated low-dimensional latent variable $z$ via an encoder function $e(\cdot)$ followed by quantization $Q(\cdot)$ to convert the continuous representation into discrete values, i.e., $z=Q(e(x))$.
  • Entropy Coding: After determining $z$, lossless entropy coding, such as Huffman coding [51] or arithmetic coding [65], is applied to produce a compressed bit sequence with length $b(z)$. Ideally, the entropy coding scheme approximates the theoretical bit rate, given by the entropy of $z$ under its marginal distribution. We assume that both the sender and receiver have access to an entropy model $P(z)$, which estimates the marginal probability distribution of $z$ and determines the expected bit length as $b(z)\approx-\log_{2}P(z)$. The goal is to ensure that $P(z)$ closely approximates the true marginal distribution $p(z)$, which is defined as $p(z)=\mathbb{E}_{x\sim p_{\text{source}}}\left[\delta(z,Q(e(x)))\right],$ where $\delta$ is the Kronecker delta function. Under this approximation, the encoding length $b(z)$ is nearly optimal, as the average code length approaches the entropy $-\log_{2}p(z)$ of $z$.
  • Reconstruction: Once the receiver has obtained $z$, it reconstructs the approximation $\hat{x}$ using a reconstruction function $r(\cdot)$, such that $\hat{x}=r(z)$.
@@
$\mathcal{L}_{\text{R-D}}=\mathbb{E}_{x\sim p_{\text{source}}}\left[\underbrace{-\log_{2}P(z)}_{\text{rate}\ (\mathcal{L}_{\text{R}})}+\underbrace{\lambda d(x,\hat{x})}_{\text{distortion}\ (\mathcal{L}_{\text{D}})}\right],$ (2)
@@ 3.2 Balanced rate-distortion optimization
Let $g(d_{t},w)=(w_{\text{R},t}\nabla\log\mathcal{L}_{\text{R},t}+w_{\text{D},t}\nabla\log\mathcal{L}_{\text{D},t})^{\top}d_{t}-\frac{1}{2}\|d_{t}\|^{2}$. The optimal direction $d_{t}^{*}$ can be obtained by setting $\frac{\partial g}{\partial d_{t}}=0$:
$\frac{\partial g}{\partial d_{t}}=0\quad\Longrightarrow\quad d_{t}^{*}=w_{\text{R},t}\nabla\log\mathcal{L}_{\text{R},t}+w_{\text{D},t}\nabla\log\mathcal{L}_{\text{D},t}.$ (8)
@@ 3.2.1 Solution 1: Gradient descent over trajectory
where $\xi_{t}\in\mathbb{R}^{2}$ represents the unconstrained softmax logits. To give more weight to recent updates, we add a decay term [75], leading to the following update for $\xi_{t}$:
@@
9:     Apply softmax for numerical stability: $w_{t}=\text{Softmax}({w}_{t})$
10:     Compute normalization constant: $c_{t}=\left(\frac{w_{\text{R},t}}{\mathcal{L}_{\text{R},t}}+\frac{w_{\text{D},t}}{\mathcal{L}_{\text{D},t}}\right)^{-1}$
11:     Compute balanced gradient: $\mathbf{d}_{t}=c_{t}\left(w_{\text{R},t}\nabla_{\theta}\log\mathcal{L}_{\text{R},t}+w_{\text{D},t}\nabla_{\theta}\log\mathcal{L}_{\text{D},t}\right)$
@@ 2. Primal Feasibility: The weights must satisfy the equality constraint,
${w}_{t}=\text{Softmax}\left(\frac{Q^{-1}\mathbf{1}}{\mathbf{1}^{\top}Q^{-1}\mathbf{1}}\right).$ (26)
This step ensures $w_{t}\in\mathbb{S}_{2}$ and reduces potential numerical issues caused by large gradient variations.
After determining the weights $w_{t}$, we apply the same renormalization constant $c_{t}$ as in Section 3.2.1 to compute a balanced gradient for updating the model parameters $\theta$, following the form in Eq. 18. Solution 2, with its analytical QP formulation, is particularly suitable for fine-tuning existing LIC models, offering a refined approach to balance rate and distortion objectives. The detailed implementation is presented in Algorithm 2.
@@ 4.4.2 Weight decay
In Solution 1, we introduce a weight decay term when updating the unconstrained softmax logits to mitigate potential non-convergence issues that arise from instantaneous gradients [75]. To evaluate the effect of weight decay, we perform a series of experiments, varying the decay coefficient $\gamma$ across values $\{0.01,0.015,0.005,0.001,0.0005,0\}$, as shown in Fig. 3(b). Notably, setting $\gamma=0$ (i.e., no weight decay) results in non-convergence issues that precluded plotting the results. Among the other tested values, $\gamma=0.001$ yields the best results, and we thus adopt $\gamma=0.001$ consistently across all values of $\lambda$ and LIC models.
@@ 5 Conclusion
Limitations and Future Work. Our approach balances the R-D optimization of LICs by introducing an additional subproblem, which moderately increases training complexity. Future work could focus on designing more efficient algorithms to mitigate this added computational cost. Recent research indicates that Tchebycheff Scalarization [39, 40] is a promising avenue, offering favorable theoretical properties along with low computational complexity. Additionally, in our solutions, we employ the softmax function to constrain weights within the probability simplex; however, projection-based methods could be also explored [33, 64].

### v2

@@ Abstract
Learned image compression (LIC) using deep learning architectures has seen significant advancements, yet standard rate-distortion (R-D) optimization often encounters imbalanced updates due to diverse gradients of the rate and distortion objectives. This imbalance can lead to suboptimal optimization, where one objective dominates, thereby reducing overall compression efficiency. To address this challenge, we reformulate R-D optimization as a multi-objective optimization (MOO) problem and introduce two balanced R-D optimization strategies that adaptively adjust gradient updates to achieve more equitable improvements in both rate and distortion. The first proposed strategy utilizes a coarse-to-fine gradient descent approach along standard R-D optimization trajectories, making it particularly suitable for training LIC models from scratch. The second proposed strategy analytically addresses the reformulated optimization as a quadratic programming problem with an equality constraint, which is ideal for fine-tuning existing models. Experimental results demonstrate that both proposed methods enhance the R-D performance of LIC models, achieving around a 2% BD-Rate reduction with acceptable additional training cost, leading to a more balanced and efficient optimization process. Code will be available at https://gitlab.com/viper-purdue/Balanced-RD.
@@ 3.1 Preliminaries: rate-distortion optimization in learned image compression
  • Encoding and Quantization: First, each data point $x$ is mapped to a de-correlated low-dimensional latent variable $\hat{z}$ via an encoder function $e(\cdot)$ followed by quantization $Q(\cdot)$ to convert the continuous representation into discrete values, i.e., $\hat{z}=Q(e(x))$.
  • Entropy Coding: After determining $\hat{z}$, lossless entropy coding, such as Huffman coding [51] or arithmetic coding [65], is applied to produce a compressed bit sequence with length $b(\hat{z})$. Ideally, the entropy coding scheme approximates the theoretical bit rate, given by the entropy of $\hat{z}$ under its marginal distribution. We assume that both the sender and receiver have access to an entropy model $P(\hat{z})$, which estimates the marginal probability distribution of $\hat{z}$ and determines the expected bit length as $b(\hat{z})\approx-\log_{2}P(\hat{z})$. The goal is to ensure that $P(\hat{z})$ closely approximates the true marginal distribution $p(\hat{z})$, which is defined as $p(\hat{z})=\mathbb{E}_{x\sim p_{\text{source}}}\left[\delta(\hat{z},Q(e(x)))\right],$ where $\delta$ is the Kronecker delta function. Under this approximation, the encoding length $b(\hat{z})$ is nearly optimal, as the average code length approaches the entropy $-\log_{2}p(\hat{z})$ of $\hat{z}$.
  • Reconstruction: Once the receiver has obtained $\hat{z}$, it reconstructs the approximation $\hat{x}$ using a reconstruction function $r(\cdot)$, such that $\hat{x}=r(\hat{z})$.
@@
$\mathcal{L}_{\text{R-D}}=\mathbb{E}_{x\sim p_{\text{source}}}\left[\underbrace{-\log_{2}P(\hat{z})}_{\text{rate}\ (\mathcal{L}_{\text{R}})}+\underbrace{\lambda d(x,\hat{x})}_{\text{distortion}\ (\mathcal{L}_{\text{D}})}\right],$ (2)
@@ 3.2 Balanced rate-distortion optimization
Let $g(d_{t},w)=(w_{\text{R},t}\nabla\log\mathcal{L}_{\text{R},t}+w_{\text{D},t}\nabla\log\mathcal{L}_{\text{D},t})^{\top}d_{t}-\frac{1}{2}\|d_{t}\|^{2}$. The optimal direction $d_{t}^{*}$ is obtained by setting:
$\small\frac{\partial g}{\partial d_{t}}=0\quad\Longrightarrow\quad d_{t}^{*}=w_{\text{R},t}\nabla\log\mathcal{L}_{\text{R},t}+w_{\text{D},t}\nabla\log\mathcal{L}_{\text{D},t}.$ (8)
@@ 3.2.1 Solution 1: Gradient descent over trajectory
where $\xi_{t}\in\mathbb{R}^{2}$ represents the unconstrained softmax logits. To give more weight to recent updates, we add a decay term [75, 42], leading to the following update for $\xi_{t}$:
@@
9:     Apply softmax for numerical stability: $\tilde{w}_{t}=\text{Softmax}({w}_{t})$
10:     Compute normalization constant: $c_{t}=\left(\frac{\tilde{w}_{\text{R},t}}{\mathcal{L}_{\text{R},t}}+\frac{\tilde{w}_{\text{D},t}}{\mathcal{L}_{\text{D},t}}\right)^{-1}$
11:     Compute balanced gradient: $\mathbf{d}_{t}=c_{t}\left(\tilde{w}_{\text{R},t}\nabla_{\theta}\log\mathcal{L}_{\text{R},t}+\tilde{w}_{\text{D},t}\nabla_{\theta}\log\mathcal{L}_{\text{D},t}\right)$
@@ 2. Primal Feasibility: The weights must satisfy the equality constraint,
$\tilde{w}_{t}=\text{Softmax}\left({w}_{t}\right)=\text{Softmax}\left(\frac{Q^{-1}\mathbf{1}}{\mathbf{1}^{\top}Q^{-1}\mathbf{1}}\right).$ (26)
This step ensures $\tilde{w}_{t}\in\mathbb{S}_{2}$ and reduces potential numerical issues caused by large gradient variations. It serves as an approximation that aligns with the optimal solution while maintaining similar performance despite slight deviations.
After determining the weights $\tilde{w}_{t}$, we apply the same renormalization constant $c_{t}$ as in Section 3.2.1 to compute a balanced gradient for updating the model parameters $\theta$, following the form in Eq. 18. Solution 2, with its analytical QP formulation, is particularly suitable for fine-tuning existing LIC models, offering a refined approach to balance rate and distortion objectives. The detailed implementation is presented in Algorithm 2.
@@ 4.4.2 Weight decay
In Solution 1, we introduce a weight decay term when updating the unconstrained softmax logits to mitigate potential non-convergence issues that arise from instantaneous gradients [75]. To evaluate the effect of weight decay, we perform a series of experiments, varying the decay coefficient $\gamma$ across values $\{0.01,0.015,0.005,0.001,0.0005,0\}$, as shown in Fig. 3(b). Notably, setting $\gamma=0$ (i.e., no weight decay) results in non-convergence issues that precluded plotting the results. Among the other tested values, $\gamma=0.001$ yields the best results, and we thus adopt $\gamma=0.001$.
@@ 5 Conclusion
Limitations and Future Work. Our approach balances the R-D optimization of LICs by introducing an additional subproblem [42, 58], which moderately increases training complexity. Future work could focus on designing more efficient algorithms to mitigate this added computational cost. Recent research indicates that Tchebycheff Scalarization [39, 40] is a promising avenue, offering favorable theoretical properties along with low computational complexity. Additionally, we employ the softmax function to constrain weights within the probability simplex; however, projection-based methods could be also explored [33, 64].